### v1

# Cryptoanalysis of a public key exchange based on circulant matrix over digital semiring

Abstract
We present a cryptanalysis of a key exchange protocol based on the digital semiring. For this purpose, we find the maximal solution of a linear system over such semiring, and use the properties of circulant matrix to demonstrate that the protocol is vulnerable. Specifically, we provide an efficient attack that recovers the shared secret key from publicly exchanged information for any instance of the digital semiring in polynomial time.
1 Introduction
The modern cryptosystems appears with the fundational work of Diffie-Hellman [13], with foundations upon the difficulty of the discrete logarithm problem in cyclic groups, especially over elliptic curves. Nevertheless, advanced in quantum computing threat the secutity of those protocols. In particular, in 1994, Shor introduced a quantum algorithm capable of efficiently solving this problem. This marked the start of the research in post-quantum cryptography, dedicated to constructing cryptographic schemes resilient to quantum attacks.
As a possible solution, Maze et al. [12] proposed a key exchange protocol using semigroup actions on sets. This papers is seen as the beginning for cryptography based on semiring, as in that paper a finite congruence simple semiring is proposed as a possible algebraic framework. These algorithm can be interpreted as a generalization of classical systems such as the Diffie-Hellman [13] and ElGamal [14] protocols, but instead of a cyclic group, they use abelian semigroups. However, the instance proposed has been shown recently to be vulnerable to cryptanalysis [15]. Despite this, their framework has inspired various cryptographic developments. For example, Kahrobaei and Koupparis [11] explored the use of non-abelian group actions, pushing the original idea into the realm of non-commutative algebra. Similarly, Gnilke and Zumbrägel [16] connected these concepts with recent progress in isogeny-based cryptography. Another extension can be found in the work of Olvera-Cravioto [17], who applied twisted group rings to design new key exchange mechanisms.
In another line of research, Grigoriev and Shpilrain investigated the use of tropical semirings as a foundation for public key cryptography, including both key exchange protocols [4, 5] and digital signature schemes [6]. Nevertheless, Kotov and Ushakov [7] introduced a heuristic algorithm known as the Kotov-Ushakov attack—which has become a standard method of cryptanalysis in tropical cryptography. Building on this, Otero et all [3] recently proposed a deterministic alternative that enhances the original attack by providing consistent results while preserving its applicability. Moreover, other schemes based on tropical algebra by Grigoriev and Shpilrain have also been shown to lack security, as demonstrated in several studies [10, 8, 9].
More recently, new directions have emerged leveraging semiring structures for practical cryptographic applications. In particular, Nassr et al. proposed a public-key encryption scheme grounded on the hardness of the two-sided digital circulant matrix action problem over a semiring initially introduced by Huang et al. [2]. In this paper, we study the security of this protocol, as well as provide a numerial example
2 Mathematical preliminaries
In this secition we will introduce some basic background on tropical semiring as well as triad semiring.
Definition 2.1.
A semiring $R$ is a non-empty set together with two operations $+$ and $\cdot$ such that $(S,+)$ is a commutative monoid, $(S,\cdot)$ is a monoid and the following distributive laws hold:
$\begin{split}a(b+c)&=ab+ac\\ (a+b)c&=ac+bc\end{split}$
We say that $(R,+,\cdot)$ is additively idempotent if $a+a=a$ for all $a\in R$.
Definition 2.2.
Let $R$ be a semiring and $(M,+)$ be a commutative semigroup with identity $0_{M}$. $M$ is a right semimodule over $R$ if there is an external operation $\cdot\ :M\times R\rightarrow M$ such that
$\begin{split}(m\cdot a)\cdot b&=m\cdot(a\cdot b),\\ m\cdot(a+b)&=m\cdot a+m\cdot b,\\ (m+n)\cdot a&=m\cdot a+n\cdot a,\\ 0_{M}\cdot a&=0_{M},\end{split}$
for all $a,b\in R$ and $m,n\in M$. We will denote $m\cdot a$ by simple concatenation $ma$.
In [1], a new additively semiring is proposed, which we will call the digits semiring.
Definition 2.3.
Let $W=\mathbb{N}\cup\{\infty\}$ and for all $g\in\mathbb{N}$, let $\delta(g)$ be the sum of all digits of $g$. The digits semiring is the semiring $(W,\oplus,\otimes)$ with
$g_{1}\oplus g_{2}=\begin{cases}g_{1}&\text{if }\delta(g_{2})<\delta(g_{1}),\\ g_{2}&\text{if }\delta(g_{2})>\delta(g_{1}),\\ \max(g_{1},g_{2})&\text{if }\delta(g_{1})=\delta(g_{2}),\end{cases}$
$g_{1}\otimes g_{2}=\begin{cases}g_{1}&\text{if }\delta(g_{1})<\delta(g_{2}),\\ g_{2}&\text{if }\delta(g_{1})>\delta(g_{2}),\\ \min(g_{1},g_{2})&\text{if }\delta(g_{1})=\delta(g_{2}),\end{cases}$
To differenciate the natural order of $\mathbb{N}$ and $W$ induced by addition, we will note $\leq_{N}$ the natural order of numbers, and $\leq_{W}$ the one given in $W$.
Over all semiring we can define the semiring of matrix with coefficients in such semiring.
Note that all additively idempotent semiring $R$ have an induced order by $a\leq_{R}b$ if and only if $a+b=b$.
Definition 2.4.
Let $R$ be a semiring. Then the set of squared matrix over $R$, $Mat_{n}(R)$, is a semiring with the usual operations
  • $(A\oplus B)_{ij}=A_{ij}\oplus B_{ij}$,
  • $(A\otimes B)_{ij}=\bigoplus_{k=1}^{n}A_{ik}\otimes B_{kj}$,
Note that if $R$ is additively idempotent, then so is $Mat_{n}(R)$
Definition 2.5.
Let $R$ be a semiring. A matrix $C\in\rm{Mat}_{n}(R)$ is called circulant if there are $c_{0},c_{1},\cdots,c_{n-1}\in R$ such that
$C=\begin{pmatrix}c_{0}&c_{n-1}&c_{n-2}&\cdots&c_{1}\\ c_{1}&c_{0}&c_{n-1}&\cdots&c_{2}\\ c_{2}&c_{1}&c_{0}&\cdots&c_{3}\\ \vdots&\vdots&\vdots&\ddots&\vdots\\ c_{n-1}&c_{n-2}&c_{n-3}&\cdots&c_{0}\end{pmatrix}$
We will denote $C$ as $C=Circ(c_{0},\cdots,c_{n-1})$
A famous result regarding the structure of circulant matrix is
Theorem 2.6.
The set $Circ_{n}(R)$ of ciruclar matrix of $n\times n$ over $R$ form a commutative subsemiring of $\rm{Mat}_{n}(R)$.
In [1], the following key exchange protocol is proposed
Protocol 2.7.
Let $W$ be the digital semiring. Alice and Bob agree matrix $M\in\rm{Mat}_{n}(W)$.
  1. 1.
    Alice chose $A_{1},A_{2}\in Circ_{n}(\mathbb{W)}$ and makes public $pk_{1}=A_{1}\otimes M\otimes A_{2}$
  2. 2.
    Bob chose $B_{1},B_{2}\in Circ_{n}(\mathbb{W)}$ and makes public $pk_{1}=B_{1}\otimes M\otimes B_{2}$
  3. 3.
    Alice computes $A_{1}\otimes pk_{2}\otimes A_{2}$ and Bob computes $B_{1}\otimes pk_{1}\otimes B_{2}$
The common key is
$A_{1}\otimes pk_{2}\otimes A_{2}=A_{1}\otimes B_{1}\otimes M\otimes B_{2}\otimes A_{2}=B_{1}\otimes A_{1}\otimes M\otimes A_{2}\otimes B_{2}=B_{1}\otimes pk_{1}\otimes B_{2}$
The security of this protocol is based on the following problems
Problem 2.8 (MAP [1]).
Let $P_{1},P_{2}\in M_{n}(G)$ be two circulant matrices, and let $U=P_{1}\otimes T\otimes P_{2}$ for an arbitrary matrix $T\in M_{n}(G)$. Given $U$ and $T$, the challenge is to obtain two circulant matrices $P_{1}$ and $P_{2}$ such that
$U=P_{1}\otimes T\otimes P_{2}.$
Problem 2.9 (CMAP [1]).
Let $P_{1},P_{2},B_{1},B_{2}\in M_{n}(W)$ be circulant matrices, $U=P_{1}\otimes T\otimes P_{2}$, and $V=B_{1}\otimes T\otimes B_{2}$ for an arbitrary matrix $T\in M_{n}(W)$. Given $U$, $V$, and $T$, the challenge is to obtain
$K=P_{1}\otimes B_{1}\otimes T\otimes B_{2}\otimes P_{2}.$
Problem 2.10 (Decision MAP).
Given $T,U\in M_{n}(W)$. Are there two circulant matrices $P_{1}$ and $P_{2}$ such that
$U=P_{1}\otimes T\otimes P_{2}$
In [1], it is shown that the solution of CMAP is identical to obtaining the matrix $K$ from the known parameters of the protocol. In addition, MAP can be transformed into the problem of solving quadratic polynomial systems on the semiring $(W,\oplus,\boxtimes)$.
Proposition 2.11 ([1]).
MAP can be transformed to the problem of solving quadratic polynomial systems on the digits semiring.
Also, in [2] the authors assert that to solve those problems are NP-problems. However, that is not true. The original paper shows that solving quadratic polynomial systems on the digits semiring is a NP-problem, nevertheless, the problems proposed are special instances of the general problem, and therefore to solve those instance do not have to bee NP problem. In particular, here we propose how to solve the CMAP problem.
Finally, we introduce the concept of maximal solution,
Definition 2.12.
Let $R$ be an additively idempotent semiring, and let $XA=Y$ be a linear system of equations. We say that $\hat{X}$ is the maximal solution of the system if and only if the two following conditions are satisfied
  1. 1.
    $\hat{X}\in R^{n}$ is a solution of the system, i.e. $\hat{X}A=Y$,
  2. 2.
    if $Z\in R^{n}$ is any other solution of the system, then $Z+\hat{X}=\hat{X}$.
This last condition is equivalent to $Z\leq\hat{X}$.
3 Cryptoanalysis of the protocol
In [3] a new method to solve linear equations over additively idempotent semiring is proposed, as well as its cryptographic applications. In [2] they assert that the cryptoanalysis on such paper can not be used against 2.7, as the private keys do not come from tropical polynomials of matrices. We will present a modification of that cryptanalysis that can be used with circulant matrix.
First, let $C_{i}=C[e_{i}]$ with $e_{i}$ the $i-$th vector of the canonical base. Then, we have that
$C[a_{1},a_{2},\cdots,a_{n}]=a_{1}C_{1}\oplus a_{2}C_{2}\cdots\oplus a_{n}C_{n}$
and therefore,
$K=C[a_{1},\cdots,a_{n}]\otimes M\otimes C[b_{1},\cdots,b_{n}]=\left(\bigoplus_{i=1}^{n}a_{i}C_{1}\right)\otimes M\otimes\left(\bigoplus_{i=1}^{n}b_{i}C_{1}\right)=\bigoplus_{i,j=1}^{n}a_{i}b_{j}C_{i}\otimes M\otimes C_{j}$ (1)
where $C_{i},M$ are known. We can observe that the solutions to this system are solutions of the following linear system over $W$
$pk_{1}=\bigoplus_{i,j=1}^{n}z_{ij}C_{i}\otimes M\otimes C_{j}$ (2)
To solve the previous linear system, we must note that
Lemma 3.1.
Let $a,b\in W$, then $a\otimes b\leq a,b$, where order is based the natural order of additively idempotent semiring.
Proof.
w.l.o.g we can assume $a\leq_{W}b$. We have that $a\leq_{W}b$ means that $\delta(a)\leq_{N}\delta(b)$ or $\delta(a)=\delta(b)$ and $a\leq_{N}b$. In both cases, $a\otimes b=a\leq_{W}b$. ∎
In [3] the following characterization of maximal solution is presented.
Theorem 3.2.
Given $(R,+,\cdot)$ an additively idempotent semiring, let $T_{i}=\{x\in R:xH_{i}+Y=Y\}$ $\forall i=1,\dots,n$. Suppose that these subsets have a maximum with respect to the order induced in $R$
$C_{i}=\max T_{i}.$
If $XH=Y$ has a solution, then $Z=(C_{1},\dots C_{n})$ is the maximal solution of the system.
Note that if $\delta(a)\leq\delta(y)$, then $x\otimes a\leq_{W}y$ for all $x\in W$, and $\delta(a)>\delta(y)$ then $\delta(x)\leq\delta(y)$. As a result, $T_{i}=W$ if $H_{i}\leq_{W}\{y_{j};\delta(y_{j})\leq\delta(h_{ij})\}$, and $x\leq_{W}\min_{W}Y$ in other case, with $\min_{W}$ the min respect to the order in $W$. Therefore,
$\max T_{i}=\begin{cases}\infty&\text{if }H_{i}\leq_{W}Y,\\ \min_{W}Y&\text{if }other\end{cases}$
As there must be a solution of the system, provided that $K=A_{1}MA_{2}$, so we can compute a solution $d_{ij}$ of $K=\oplus_{i,j=1}^{n}z_{ij}C_{i}\otimes M\otimes C_{j}$. As it is done in [3], we have that
Theorem 3.3.
Let $d_{ij}$ be the maximal solution of $K=A_{1}MA_{2}=\oplus_{i,j=1}^{n}d_{ij}C_{i}\otimes M\otimes C_{j}$, and let $F[X]=\oplus_{i,j=1}^{n}d_{ij}C_{i}\otimes X\otimes C_{j}$. Then $F[M]=pk_{1}$ and $F[pk_{2}]=K_{ab}$
Proof.
$F[M]=pk_{1}$ is true by definition of $d_{ij}$. For the second one, we have
$F[pk_{2}]=$ $\bigoplus_{i,j=1}^{n}d_{ij}C_{i}pk_{2}C_{j}=\bigoplus_{i,j=1}^{n}d_{ij}C_{i}B_{1}MB_{1}C_{j}=$
$=$ $\bigoplus_{i,j=1}^{n}d_{ij}B_{1}C_{i}MC_{j}B_{2}=B_{1}\left(\bigoplus_{i,j=1}^{n}d_{ij}C_{i}MC_{j}\right)B_{2}=$
$=$ $B_{1}pk_{1}B_{2}=K$
Finally, to compute $F[X]$ needs the computation of $n^{2}$ $\delta$ and $n^{2}$ comparations in $\mathbb{N}$. If the protocol is runned with coefficients bounded by $J$, the computational cost of $\delta$ is $\log_{10}(J)$ So the computational cost of computing $F[X]$ is $o(\log_{10}(J)n^{2})$ opeations in $\mathbb{N}$
4 Conclusions
We have cryptoanalyced the public key exchange proposed in [1] and [2] based on digital semiring. To do so, we have used the original ideas of [3] to the special case of digital sum, find a method to obtain the maximal solution of a linear system over such semiring. In addition, we have used the properties of circulant matrix to obtain the private key with the use of the public parameters, showing that the key exchange protocol is vulnerable and therefore not secure.
Acknowledgments
This research was supported by the Spanish Ministry of Science, Innovation and Universities under the FPU 2023 grant program.